# The sup-norm problem for newforms of large level on $\operatorname{PGL}(n)$

### v1

@@ 7.1. The Fourier bound
$\phi(z)\ll_{\Omega,\mu\varepsilon}N^{-1/4+1/4n+\varepsilon}.$
@@
We use the bound given in Theorem 3 of [BHM20], making the necessary adjustments from the level $1$ results to level $N$. The proof is very similar, so we refer to [BHM20] for more details and mostly remark on what changes need to be made. Note first that the method of proof involves the Whittaker expansion [BHM20, (46)]. An automorphic form for the group $\Gamma_{0}(N)$ enjoys the same type of Whittaker expansion, since $\operatorname{SL}_{n-1}(\mathbb{Z})$ embeds in the upper left $(n-1)\times(n-1)$ block of $\Gamma_{0}(N)$, so that one can follow the same arguments given in, for instance, [Gol06, Theorem 5.3.2] in level $1$. To follow the arguments in [BHM20] further, we normalise $\phi$ arithmetically, so that the first coefficient in the expansion is $1$. Next, the bound [BHM20, (49)] for $L(1+\varepsilon,\pi\times\tilde{\pi})$ holds similarly, with an additional $N^{\varepsilon}$ on the right-hand side. Here, we let $\pi$ be the automorphic representation generated by $\phi$. Finally, to account for the factor between arithmetically normalised forms and $L^{2}$-normalised forms, we note the display before [BHM20, (66)]. More precisely, if we assume $\phi$ to be arithmetically normalised, as in [BHM20, (46)], then standard Rankin-Selberg theory shows that
$\left\lVert\phi\right\rVert^{2}\asymp_{\mu}\operatorname{vol}(\Gamma_{0}(N)\backslash\mathbb{H})\cdot\operatorname{res}_{s=1}L(s,\pi\times\tilde{\pi}).$
By [Bru06, Theorem 3], as in the two displays after (43) in [Lap13, Appendix], we can use the lower bound
@@
where $C(\pi\times\tilde{\pi})=C(\pi\times\tilde{\pi},0)$ is the analytic conductor of $L(s,\pi\times\tilde{\pi})$. We have $C(\pi)\asymp_{\mu}N$ and by [BH97, (2)] the bound
$C(\pi\times\tilde{\pi})\ll N^{n\cdot 1+n\cdot 1-1}$
holds. It is easy to compute that $\operatorname{vol}(\Gamma_{0}(N)\backslash\mathbb{H})\asymp N^{n-1}$. Therefore,
$\left\lVert\phi\right\rVert^{2}\gg_{\mu}N^{n-1}\cdot N^{(2n-1)(-1/2+1/2n)-\varepsilon}=N^{1/2-1/2n-\varepsilon}.$
Going back to $\phi$ being $L^{2}$-normalised by putting together the bound above and [BHM20, (49)] with the indicated adjustments, we deduce the claim. ∎
@@
Working more precisely, one could prove that $C(\pi\times\tilde{\pi})\asymp N^{2n-2}$ and improve the exponent in the bound above. This is not necessary for this paper and we do not prove the claim. In fact, one expects that $\operatorname{res}_{s=1}L(s,\pi\times\tilde{\pi})\gg N^{\varepsilon}$. This is proven by Hoffstein-Lockhart in the case $n=2$ and for this reason we have
@@
for $z\in\Omega$ as in [HT12, Lemma 4], for example.
7.2. Finishing the proof
First assume Hypothesis (1.1). Proposition 3.3 and Proposition 6.2 together with the prime number theorem imply that
under the assumptions on $z$ specified in Proposition 6.2. Optimising the size of $L$, we choose $L=N^{1/n}$. In this case, we have
$\phi(z)\ll L^{-1/4+\varepsilon}\ll N^{-1/4n+\varepsilon}.$
The Fourier bound, Proposition 7.1, certainly implies the same bound
$\phi(z)\ll N^{-1/4n+\varepsilon}$
for $n\geq 2$.
These bounds are valid on the subsets of $\mathbb{H}$ given in Proposition 5.2. As remarked at the beginning of Section 5, these now extend to the whole domain $\Omega_{N}$, and the proof is finished.
for $L\ll N^{1/n^{2}-\varepsilon}$. Maximising $L$, we get
$\phi(z)\ll L^{-1/2+\varepsilon}\ll N^{-1/2n^{2}+\varepsilon}.$
@@ Acknowledgements
I am very grateful to Valentin Blomer for suggesting the problem and for his generous and valuable support. I also thank Edgar Assing, Gergely Harcos, and Aurel Page for useful discussions on topics related to this paper.

### v2

@@ 8. Appendix: The Fourier bound
$\phi(z)\ll_{\Omega,\mu\varepsilon}N^{-1/2+1/2n+\varepsilon}.$
@@
We use the bound given in Theorem 3 of [4], making the necessary adjustments from the level $1$ results to level $N$. The proof is very similar, so we refer to [4] for more details and mostly remark on what changes need to be made.
Note first that the method of proof involves the Whittaker expansion [4, (46)]. An automorphic form for the group $\Gamma_{0}(N)$ enjoys the same type of Whittaker expansion, since $\operatorname{SL}_{n-1}(\mathbb{Z})$ embeds in the upper left $(n-1)\times(n-1)$ block of $\Gamma_{0}(N)$, so that one can follow the same arguments given in, for instance, [20, Theorem 5.3.2] in level $1$. To follow the arguments in [4] further, we normalise $\phi$ arithmetically, so that the first coefficient in the expansion is $1$.
[…]
By [14, Theorem 3], as in the two displays after (43) in [30, Appendix], we can use the lower bound
@@
[…]
$C(\pi\times\tilde{\pi})\ll N^{n\cdot 1+n\cdot 1-2}$
holds.
It is easy to compute that $\operatorname{vol}(\Gamma_{0}(N)\backslash\mathbb{H})\asymp N^{n-1}$. Therefore,
$\left\lVert\phi\right\rVert^{2}\gg_{\mu}N^{n-1}\cdot N^{(2n-2)(-1/2+1/2n)-\varepsilon}=N^{1-1/n-\varepsilon}.$
Going back to $\phi$ being $L^{2}$-normalised by putting together the bound above and [4, (49)] with the indicated adjustments, we deduce the claim. ∎
@@
In fact, one expects that $\operatorname{res}_{s=1}L(s,\pi\times\tilde{\pi})\gg N^{-\varepsilon}$. This is proven by Hoffstein-Lockhart in the case $n=2$ and for this reason we have
@@
for $z\in\Omega$ as in [21, Lemma 4], for example. This is the best-possible bound on the global sup-norm in this case.
Remark 8.3.
More precisely, for $n>2$, Blomer–Harcos–Maga [4, (63)] prove a bound of the form
$\phi(z)\ll\left(\prod_{i=1}^{n-1}y_{i}^{(n-i)(n-i-1)/2}\right)^{-1}=y_{1}^{-(n-1)(n-2)/2}\cdots y_{n-2}^{-1}\cdot y_{n-1}^{0},$
assuming that $y_{1},\ldots,y_{n-1}\gg 1$, e.g. $z$ is in a standard fundamental domain for $\operatorname{SL}_{n}(\mathbb{Z})$. Even if this would hold for smaller $y_{1}$ and assuming a Hoffstein-Lockhart-type bound, one would obtain a sub-baseline bound only for $y_{1}\gg N^{-2/(n-2)+\varepsilon}$ and $y_{i}\gg 1$ for all $i\geq 2$. This is still very far from covering the region $C_{N}$.
However, the form of the bound, given through a product of $y$-coordinates, does suggest decay as soon as at least one of the coordinates is significantly larger than in the bulk. This is strengthened in [4, (6.5)] using the dual form of $\phi$, where all $y$-coordinates appear with negative exponents. However, when $N>1$, the latter has a more complicated Whittaker expansion, which deserves to be investigated.
@@ Acknowledgements
I am very grateful to Valentin Blomer for suggesting the problem and for his generous and valuable support. Part of this paper was worked out while I was a visiting student at EPFL, and I am grateful to Philippe Michel and his group for great working conditions. I also thank Edgar Assing, Farrell Brumley, Gergely Harcos, and Aurel Page for useful discussions on topics related to this paper, and the referee for improving readability and correctness.